# Extracting the Structure of Press Releases for Predicting Earnings Announcement Returns

### v1

@@ Abstract.
We examine how textual features in earnings press releases predict stock returns on earnings announcement days. Using over 138,000 press releases from 2005 to 2023, we compare traditional bag-of-words and BERT-based embeddings. We find that press release content (soft information) is as informative as earnings surprise (hard information), with FinBERT yielding the highest predictive power. Combining models enhances explanatory strength and interpretability of the content of press releases. Stock prices fully reflect the content of press releases at market open. If press releases are leaked, it offers predictive advantage. Topic analysis reveals self-serving bias in managerial narratives. Our framework supports real-time return prediction through the integration of online learning, provides interpretability and reveals the nuanced role of language in price formation.
@@ 1. Introduction
Earnings announcements are some of the most important anticipated scheduled announcements by corporations in financial markets, often triggering immediate and substantial price adjustments. While research has traditionally focused on the “hard” components of earnings releases—namely, the numerical earnings surprises that deviate from analyst expectations—a growing body of research show that substantial market-relevant information is also embedded in the “soft” content of these announcements. In particular, earnings press releases—which are the company-issued quarterly performance statements disseminated as the initial source of public news are a major vehicle through which both hard and soft information reaches investors. These texts often include management’s interpretation of performance, strategic positioning, and forward-looking commentary, all of which can influence investor expectations about the firm’s future prospects. Despite their importance, earnings press releases have received relatively limited attention in the financial literature, especially when compared to conference calls. This is surprising given that press releases are usually released first, are more widely disseminated, and are written in more structured and accessible language than conference call transcripts. Moreover, press releases are the documents that were most often targeted in high-profile cybersecurity breaches—such as the earnings press release hacking scandal documented by Akey et al. (2022)—underscoring their value in shaping market expectations.
This paper investigates whether and how the textual content of earnings press releases predicts earnings announcement date stock return responses. Using a comprehensive dataset of over 138,000 press releases issued between 2005 and 2023, we evaluate the predictive power of soft information using a variety of natural language processing (NLP) techniques. These include traditional bag-of-words models, specifically Latent Dirichlet Allocation (LDA), and contextual embedding models from the BERT family, including FinBERT—fine-tuned specifically for financial text. We construct real-time return forecasts using a rolling window approach that avoids look-ahead bias, and we assess explanatory power of soft information embedded in press releases while controlling for earnings surprises using both regression and Shapley values.
@@ 2. Related Work
A number of studies examine the content of earnings conference calls and how it relates to stock returns (e.g., Price et al., 2012; Meursault et al., 2023; Garcia et al., 2023; Cao et al., 2024). However, only a few studies examined the content of earnings press releases (e.g., Huang et al., 2014; Akey et al., 2022). Press releases are important because they represent the main piece of news that is issued by corporations at the time of the announcement that initiates first price adjustment to earnings news. On the methodology front, Cao et al. (2024) employs multi-model analysis of conference call transcripts and audio to predict stock volatility, but their sample is limited to 500 firms in 2017. In contrast, we analyze press releases for over 6,000 unique stocks from 2006 to 2023, with time-varying return predictions and better interpretability. We further differentiate from Cao et al. (2024) since we examine the combination of both hard and soft information in how it impacts first moments, i.e,. stock returns, and how fast the information is impounded into stock prices.
@@ 3. Data
Our sample period begins January 2005 and ends in December 2023. We obtain earnings dates and times, analysts’ forecasts, and announced EPS from LSEG IBES Academic. We retrieve daily stocks returns and open and close prices from CRSP Daily Stock Files and intraday quotes from NYSE Trade and Quote (TAQ). Following (Martineau, 2022), we only select earnings announcements for which there is analyst following so that we compute earnings surprises (hard information) as
@@
where $EPS_{c,\tau}$ is firm $c$’s earnings per share of quarterly earnings announcement announced on date $\tau$, and $E_{\tau-1}[EPS_{c,\tau}]$ is the expected earnings per share measured as the concensus analyst forecasts. The consensus analyst forecast is the median of all latest analyst forecasts issued over the 90 days before the earnings announcement date. We scale the surprise by the firm’s stock price $P_{c,\tau-5}$ five trading days before the announcement.
We retrieve earnings press releases, in HTML format, published as attachments to 8-K fillings from SEC Edgar. We retrieve 158,797 earnings press releases for 6,543 unique firms (PERMNOs in CRSP). After merging with CRSP daily stock returns, we have a total of 140,425 unique stock-earnings announcement observations. After preprocessing steps of earnings press releases (removing articles that are too short or too long), we have 138,676 stock-earnings announcement observations. Since 2005 is our initial training sample, our main analysis dataset contains 134,354 press releases from 2006 to 2023.
@@
Despite these efforts, variability in press release formats and the unstructured nature of text data introduce residual noise. Futher, for bag-of-words vectorization, we apply standard text preprocessing: unescape HTML characters; remove URLs, email addresses, numbers, and special symbols; normalize whitespace and newlines; lemmatize tokens; and remove stop words and repeated patterns (e.g., “year year”, “month month”).
@@ 4.1. Bag-of-Words
We use the BKMX taxanomy from (Bybee et al., 2024) as benchmark, which defines 180 topics across 3,592 phrases, spanning economics and politics. Each article is encoded as a 180-dimensional topic vector, and topics are grouped into hierarchical metatopics. In our experiments, we chose 11 metatopics of interest. However, the BKMX model, trained on Wall Street Journal articles from 1984-2017, may not generalize well to press releases. It includes forward-looking tokens (e.g. “iphone” in Mobile device, “twitter” in Internet), while missing newer terms (e.g. “president Biden”). Ambiguities also arise. Similar tokens may appear in multiple topics (e.g. “cit” appears in both Buffett and Major concerns).
@@ 4.2. BERT Models
Bi-directional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is a transformer-based language model pre-trained on large text corpora and adaptable to downstream tasks through fine-tuning. We use several pre-trained models from the BERT family, BERT-base-uncased (Devlin et al., 2019), MPNET (Song et al., 2020), and FinBERT (Araci, 2019), to extract contextual embeddings for each article. Specifically, we represent each article as a 768-dimensional vector by applying mean pooling to the last hidden states of the first 512 tokens. FinBERT is a domain-specific variant fine-tuned for financial sentiment analysis using the Reuters TRC2 corpus and Financial PhraseBank. While standard BERT models rely on masked language modeling, which captures token-level context but ignores inter-token dependencies, MPNET improves on this by incorporating permuted language modeling to better learn both positional and contextual relationships. Cosine similarities between BERT and FinBERT embeddings are moderately high, with an average of $0.72\pm 0.03$, ranging from 0.6 to 0.8, indicating alignment in semantics. In contrast, embeddings from BERT and MPNET are mostly orthogonal, with a mean similarity of $0.01\pm 0.03$, suggesting they capture distinct features. This diversity in representations indicates that combining models may help capture complementary aspects of the text and improve downstream performance.
@@ 5.1. Return Score Construction
where $I_{j}$ is the mean absolute SHAP value for feature $j\in X$ across $N$ samples, representing the expected difference in the model outcomes $y$ when $j$ is added to a subset of features $S\subset X\setminus\left\{j\right\}$, and $\hat{I}_{j}$ is normalized across all $n$ features. Table 1 summarizes the results. Column (1) reports that in an univariate analysis, $Surprise$ is positive and statistically significant at the 1%. A one standard deviation (0.02) increase in the earnings surprise is associated with a 1.6% (0.02 $\times$ 0.81 $\approx$ 1.6%) increase in returns. The $R^{2}$ indicates that earnings surprise explains 3.3% of the variation in announcement date returns. Columns (2)-(6) controls for the soft information from press releases estimated from different models. Across all model specifications, Soft is positive and statistically significant at the 1%, increasing the $R^{2}$ up to 4.4%. The soft model estimated through FinBERT corresponds to the largest increase in $R^{2}$ relative to the univariate analysis in Column (1). Column (6) reports that a one standard deviation increase in Soft (0.016) is associated with an increase of 1% in returns (0.016 $\times$ 0.60 $\approx$ 1%).
@@
Next, we evaluate the predictive power of soft information in predicting earnings announcement date returns when including all the constructed soft variables into the following regression:
@@
To further emphasize the importance of capturing soft information embedded in press releases to predict the price revelation on earnings announcement days, Figure 2 shows the average return heatmap, sorted by quarterly quintiles of $\text{Soft}^{Mean}$ and realized earnings surprise. We define an aggregated textual signal, $\text{Soft}^{Mean}$, as the average of soft variables from BKMX, oLDA, BERT, MPNET, and FinBERT, to provide a unified measure of sentiment. The heatmap shows that the average return is driven by the composite signal, with the highest returns of 4.58% in the upper-right corner and the lowest returns of -5.06% in the bottom-left corner. The heatmap emphasize that a trader that only processes the earning surprise (hard information) will fail to capture much of the variation in return that is driven by the soft information.
@@ 5.2. Topic Analysis
For the BKMX method, we rely on the original taxonomy, from which we chose 11 metatopics. In the oLDA approach, we use the 50 topics identified by the model and group them into 9 metatopics using GPT-4o. For BERT-based models, we start with extracting influential tokens in a manner analogous to bag-of-words techniques. Since Lasso regression identifies which of the 768 BERT features are positively or negatively associated with returns, we map these weights back to tokens. Let $E\in\mathbb{R}^{512\times 768}$ be the BERT embedding matrix for the first 512 tokens in an article, and $w\in\mathbb{R}^{768}$ be the Lasso weight vector. The importance score vector $IS\in\mathbb{R}^{512}$ for each token is computed as $IS=E\cdot w$. We then filter out uninformative tokens, including: (1) wordpiece tokens starting with # [^2] We could instead merge wordpiece tokens into their parent tokens, but the current extracted tokens are already informative.; (2) punctuation, numeric and single character tokens; and (3) BERT-specific tokens [CLS], [SEP], [PAD], [MASK], [UNK]. From the remaining tokens, we identify the top 5 positive and negative tokens per article, then group them into topics. This results in approximately 13,000 unique tokens from 2006 to 2023. These tokens are divided into chunks of 500 and labeled using GPT-4o, mapped to the oLDA metatopics due to its greater relevance to earnings-related content. Tokens that GPT-4o fails to categorize are assigned to a generic “Other” category and excluded from further analysis. In total, GPT-4o successfully categorized 3,219 tokens for BERT, 3,065 for MPNET, and 2,382 for FinBERT, representing around 20% of all extracted tokens. Token counts per topic over all articles are then normalized to sum to 100% within each model.
@@ 6.1. Market Efficiency
    Strategy 1 ($j=1$): Enter at ask (bid) if signal is positive (negative), and exit at the close price. $\text{Ret}_{\tau}^{OC,1}=\frac{p_{\tau}}{p_{\tau}^{o,k}}-1$, where $p_{\tau}$ is close price on day $\tau$, $p_{\tau}^{o,k}$ is the best ask/bid price at 9:45 a.m on day $\tau$, and $k\in\left\{\text{Ask},\text{Bid}\right\}$.
@@
    Strategy 2 ($j=2$): Enter and exit using ask/bid quotes, capturing transaction costs. For instance, a long position buys at ask and sells at bid. $\text{Ret}_{\tau}^{OC,2}$ is computed in a similar fashion, with $p_{\tau}$ the close price replaced by ask/bid price at close.
@@
In (Akey et al., 2022) the authors document a case where hackers infiltrated US-based newswire services–notably PR Newswire, Business Wire, and Marketwired–and stole thousands of yet-to-be-published earnings press releases. The hackers sold this information to traders, who executed front-running trades just before the official release of the earnings announcements. This hacking operation occurred over several years (roughly 2010 to 2015) and the traders pocketed more than 100 million US dollars. Eventually, the SEC and DOJ charged more than 30 individuals in one of the largest securities fraud cases of its kind. The incident highlight that press releases are a critical node in the information dissemination chain, and any breach undermines investor confidence, market efficiency, and regulatory fairness.
@@
Our method and results so far as demonstrated how important are earnings surprises and press releases in explaining announcement dates returns, but if one ranks stocks based on expected returns and evaluate the ranking relative to the ground truth (i.e., realized returns), what fraction can we correctly identify?
@@ Acknowledgements.
This research was carried out at Rotman School of Management, FinHub Financial Innovation Lab, University of Toronto. We gratefully acknowledges financial support from the Social Sciences and Humanities Research Council (SSHRC) of Canada grant number 435-2022-0745.

### v2

@@ Abstract.
We examine how textual features in earnings press releases predict stock returns on earnings announcement days. Using over 138,000 press releases from 2005 to 2023, we compare traditional bag-of-words and BERT-based embeddings. We find that press release content (soft information) is as informative as earnings surprise (hard information), with FinBERT yielding the highest predictive power. Combining models enhances the explanatory strength and interpretability of the content of press releases. Stock prices fully reflect the content of press releases at market open. If press releases are leaked, it offers predictive advantage. Topic analysis reveals self-serving bias in managerial narratives. Our framework supports real-time return prediction through the integration of online learning, provides interpretability and reveals the nuanced role of language in price formation.
@@ 1. Introduction
Earnings announcements are some of the most important anticipated scheduled announcements by corporations in financial markets, often triggering immediate and substantial price adjustments. While research has traditionally focused on the “hard” components of earnings releases—namely, the numerical earnings surprises that deviate from analyst expectations—a growing body of research shows that substantial market-relevant information is also embedded in the “soft” content of these announcements. In particular, earnings press releases—which are the company-issued quarterly performance statements disseminated as the initial source of public news are a major vehicle through which both hard and soft information reaches investors. These texts often include management’s interpretation of performance, strategic positioning, and forward-looking commentary, all of which can influence investor expectations about the firm’s future prospects. Despite their importance, earnings press releases have received relatively limited attention in the financial literature, especially when compared to conference calls. This is surprising given that press releases are usually released first, are more widely disseminated, and are written in more structured and accessible language than conference call transcripts. Moreover, press releases are the documents that were most often targeted in high-profile cybersecurity breaches—such as the earnings press release hacking scandal documented by Akey et al. (2022)—underscoring their value in shaping market expectations.
This paper investigates whether and how the textual content of earnings press releases predicts earnings announcement date stock return responses. Using a comprehensive dataset of over 138,000 press releases issued between 2005 and 2023, we evaluate the predictive power of soft information using a variety of natural language processing (NLP) techniques. These include traditional bag-of-words models, specifically Latent Dirichlet Allocation (LDA), and contextual embedding models from the BERT family, including FinBERT—fine-tuned specifically for financial text. We construct real-time return forecasts using a rolling window approach that avoids look-ahead bias, and we assess the explanatory power of soft information embedded in press releases while controlling for earnings surprises using both regression and Shapley values.
@@ 2. Related Work
A number of studies examine the content of earnings conference calls and how it relates to stock returns (e.g., Price et al., 2012; Meursault et al., 2023; Garcia et al., 2023; Cao et al., 2024). However, only a few studies examined the content of earnings press releases (e.g., Huang et al., 2014; Akey et al., 2022). Press releases are important because they represent the main piece of news that is issued by corporations at the time of the announcement that initiates the first price adjustment to earnings news. On the methodology front, Cao et al. (2024) employ multi-model analysis of conference call transcripts and audio to predict stock volatility, but their sample is limited to 500 firms in 2017. In contrast, we analyze press releases for over 6,000 unique stocks from 2006 to 2023, with time-varying return predictions and better interpretability. We further differentiate from Cao et al. (2024) since we examine the combination of both hard and soft information in how it impacts first moments, i.e,. stock returns, and how fast the information is impounded into stock prices.
@@ 3. Data
Our sample period begins in January 2005 and ends in December 2023. We obtain earnings dates and times, analysts’ forecasts, and announced EPS from LSEG IBES Academic. We retrieve daily stock returns and open and closing prices from CRSP Daily Stock Files and intraday quotes from NYSE Trade and Quote (TAQ). Following (Martineau, 2022), we only select earnings announcements for which there is analyst following so that we compute earnings surprises (hard information) as
@@
where $EPS_{c,\tau}$ is firm $c$’s earnings per share of quarterly earnings announcement announced on date $\tau$, and $E_{\tau-1}[EPS_{c,\tau}]$ is the expected earnings per share measured as the consensus analyst forecasts. The consensus analyst forecast is the median of all latest analyst forecasts issued over the 90 days before the earnings announcement date. We scale the surprise by the firm’s stock price $P_{c,\tau-5}$ five trading days before the announcement.
We retrieve earnings press releases, in HTML format, published as attachments to 8-K filings from SEC Edgar. We retrieve 158,797 earnings press releases for 6,543 unique firms (PERMNOs in CRSP). After merging with CRSP daily stock returns, we have a total of 140,425 unique stock-earnings announcement observations. After preprocessing steps of earnings press releases (removing articles that are too short or too long), we have 138,676 stock-earnings announcement observations. Since 2005 is our initial training sample, our main analysis dataset contains 134,354 press releases from 2006 to 2023.
@@
Despite these efforts, variability in press release formats and the unstructured nature of text data introduce residual noise. Further, for bag-of-words vectorization, we apply standard text preprocessing: unescape HTML characters; remove URLs, email addresses, numbers, and special symbols; normalize whitespace and newlines; lemmatize tokens; and remove stop words and repeated patterns (e.g., “year year”, “month month”).
@@ 4.1. Bag-of-Words
We use the BKMX taxonomy from (Bybee et al., 2024) as benchmark, which defines 180 topics across 3,592 phrases, spanning economics and politics. Each article is encoded as a 180-dimensional topic vector, and topics are grouped into hierarchical metatopics. In our experiments, we chose 11 metatopics of interest. However, the BKMX model, trained on Wall Street Journal articles from 1984-2017, may not generalize well to press releases. It includes forward-looking tokens (e.g. “iphone” in Mobile device, “twitter” in Internet), while missing newer terms (e.g. “president Biden”). Ambiguities also arise. Similar tokens may appear in multiple topics (e.g. “cit” appears in both Buffett and Major concerns).
@@ 4.2. BERT Models
Bi-directional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is a transformer-based language model pre-trained on large text corpora and adaptable to downstream tasks through fine-tuning. We use several pre-trained models from the BERT family, BERT-base-uncased (Devlin et al., 2019), MPNET (Song et al., 2020), and FinBERT (Araci, 2019), to extract contextual embeddings for each article. [^2] To minimize potential forward-looking bias in the training data, we focus on models developed prior to 2020. Specifically, we represent each article as a 768-dimensional vector by applying mean pooling to the last hidden states of the first 512 tokens. FinBERT is a domain-specific variant fine-tuned for financial sentiment analysis using the Reuters TRC2 corpus and Financial PhraseBank. While standard BERT models rely on masked language modeling, which captures token-level context but ignores inter-token dependencies, MPNET improves on this by incorporating permuted language modeling to better learn both positional and contextual relationships. Cosine similarities between BERT and FinBERT embeddings are moderately high, with an average of $0.72\pm 0.03$, ranging from 0.6 to 0.8, indicating alignment in semantics. In contrast, embeddings from BERT and MPNET are mostly orthogonal, with a mean similarity of $0.01\pm 0.03$, suggesting they capture distinct features. This diversity in representations indicates that combining models may help capture complementary aspects of the text and improve downstream performance.
@@ 5.1. Return Score Construction
where $I_{j}$ is the mean absolute SHAP value for feature $j\in X$ across $N$ samples, representing the expected difference in the model outcomes $y$ when $j$ is added to a subset of features $S\subset X\setminus\left\{j\right\}$, and $\hat{I}_{j}$ is normalized across all $n$ features. Table 1 summarizes the results. Column (1) reports that in an univariate analysis, $Surprise$ is positive and statistically significant at the 1%. A one standard deviation (0.02) increase in the earnings surprise is associated with a 1.6% (0.02 $\times$ 0.81 $\approx$ 1.6%) increase in returns. The $R^{2}$ indicates that earnings surprise explains 3.3% of the variation in announcement date returns. Columns (2)-(6) control for the soft information from press releases estimated from different models. Across all model specifications, Soft is positive and statistically significant at the 1%, increasing the $R^{2}$ up to 4.4%. [^3] Surprise tends to bias the $R^{2}$ downward. It can be improved by using deciles of surprise instead of the actual level. The range of $R^{2}$ values we obtain is consistent with the literature on returns to surprise (Bird et al., 2019; Martineau, 2022; Akey et al., 2022). The soft model estimated through FinBERT corresponds to the largest increase in $R^{2}$ relative to the univariate analysis in Column (1). Column (6) reports that a one standard deviation increase in Soft (0.016) is associated with an increase of 1% in returns (0.016 $\times$ 0.60 $\approx$ 1%).
@@
Next, we evaluate the predictive power of soft information in predicting earnings announcement date returns when including all the constructed soft variables in the following regression:
@@
To further emphasize the importance of capturing soft information embedded in press releases to predict the price revelation on earnings announcement days, Figure 2 shows the average return heatmap, sorted by quarterly quintiles of $\text{Soft}^{Mean}$ and realized earnings surprise. We define an aggregated textual signal, $\text{Soft}^{Mean}$, as the average of soft variables from BKMX, oLDA, BERT, MPNET, and FinBERT, to provide a unified measure of sentiment. The heatmap shows that the average return is driven by the composite signal, with the highest returns of 4.58% in the upper-right corner and the lowest returns of -5.06% in the bottom-left corner. The heatmap emphasizes that a trader that only processes the earning surprise (hard information) will fail to capture much of the variation in return that is driven by the soft information.
@@ 5.2. Topic Analysis
For the BKMX method, we rely on the original taxonomy, from which we chose 11 metatopics. In the oLDA approach, we use the 50 topics identified by the model and group them into 9 metatopics using GPT-4o. For BERT-based models, we start with extracting influential tokens in a manner analogous to bag-of-words techniques. Since Lasso regression identifies which of the 768 BERT features are positively or negatively associated with returns, we map these weights back to tokens. Let $E\in\mathbb{R}^{512\times 768}$ be the BERT embedding matrix for the first 512 tokens in an article, and $w\in\mathbb{R}^{768}$ be the Lasso weight vector. The importance score vector $IS\in\mathbb{R}^{512}$ for each token is computed as $IS=E\cdot w$. We then filter out uninformative tokens, including: (1) wordpiece tokens starting with # [^4] We could instead merge wordpiece tokens into their parent tokens, but the current extracted tokens are already informative.; (2) punctuation, numeric and single-character tokens; and (3) BERT-specific tokens [CLS], [SEP], [PAD], [MASK], [UNK]. From the remaining tokens, we identify the top 5 positive and negative tokens per article, then group them into topics. This results in approximately 13,000 unique tokens from 2006 to 2023. These tokens are divided into chunks of 500 and labeled using GPT-4o, mapped to the oLDA metatopics due to its greater relevance to earnings-related content. Tokens that GPT-4o fails to categorize are assigned to a generic “Other” category and excluded from further analysis. In total, GPT-4o successfully categorized 3,219 tokens for BERT, 3,065 for MPNET, and 2,382 for FinBERT, representing around 20% of all extracted tokens. Token counts per topic over all articles are then normalized to sum to 100% within each model.
@@ 6.1. Market Efficiency
    Strategy 1 ($j=1$): Enter at ask (bid) if the signal is positive (negative), and exit at the closing price. $\text{Ret}_{\tau}^{OC,1}=\frac{p_{\tau}}{p_{\tau}^{o,k}}-1$, where $p_{\tau}$ is the closing price on day $\tau$, $p_{\tau}^{o,k}$ is the best ask/bid price at 9:45 a.m on day $\tau$, and $k\in\left\{\text{Ask},\text{Bid}\right\}$.
@@
    Strategy 2 ($j=2$): Enter and exit using ask/bid quotes, capturing transaction costs. For instance, a long position buys at ask and sells at bid. $\text{Ret}_{\tau}^{OC,2}$ is computed in a similar fashion, with $p_{\tau}$ the closing price replaced by ask/bid price at close.
@@
In (Akey et al., 2022) the authors document a case where hackers infiltrated US-based newswire services–notably PR Newswire, Business Wire, and Marketwired–and stole thousands of yet-to-be-published earnings press releases. The hackers sold this information to traders, who executed front-running trades just before the official release of the earnings announcements. This hacking operation occurred over several years (roughly 2010 to 2015) and the traders pocketed more than 100 million US dollars. Eventually, the SEC and DOJ charged more than 30 individuals in one of the largest securities fraud cases of its kind. The incident highlights that press releases are a critical node in the information dissemination chain, and any breach undermines investor confidence, market efficiency, and regulatory fairness.
@@
Our method and results so far have demonstrated how important are earnings surprises and press releases in explaining announcement date returns, but if one ranks stocks based on expected returns and evaluate the ranking relative to the ground truth (i.e., realized returns), what fraction can we correctly identify?
@@ Acknowledgements.
This research was carried out at Rotman School of Management, FinHub Financial Innovation Lab, University of Toronto. We gratefully acknowledge financial support from the Social Sciences and Humanities Research Council (SSHRC) of Canada grant number 435-2022-0745.